# Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation

### v1

@@ 4.2 Comparison with State-of-the-Art Methods
PASCAL VOC.   Three main experiments are conducted on the Original, Blended, and Priority protocols, which are divided according to the criteria for selecting labeled images. In Table I, the proposed method outperforms state-of-the-art methods in almost all partitions, even with a single trainable network and a lower resolution. In particular, PrevMatch significantly improves performance in settings with fewer labels (92 and 183). Compared to Diverse Co-T, PrevMatch exhibits comparable performance in settings with more than 366 labels, but it operates on a single trainable network, indicating that the proposed method is efficient with respect to training costs. For Blended and Priority protocols reported in Tables II and III, the proposed method consistently outperforms existing methods.
Cityscapes.   Experimental results for the Cityscapes validation set are listed in Table IV. Under four label partitions using ResNet-50 and ResNet-101 backbones, the proposed method outperforms previous methods in six out of eight cases. Similar to the results reported in Table 1, PrevMatch consistently improves performance in setups with fewer labels. However, certain methods (e.g., AugSeg, AEL, and Dual Teacher) perform better than PrevMatch at the 1/2 label setups. This can be attributed to the advanced data augmentation techniques used in these methods. In other words, these methods could achieve better performance because the 1/2 label partition increasingly resembles a supervised learning setting, where augmentation techniques play a crucial role. Thus, we intend to explore the efficacy of the advanced augmentations on the proposed method in future work.
TABLE IV: Comparison with state-of-the-art methods on the Cityscapes dataset. All methods are trained using ResNet-50/101 and DeepLabV3+.
Cityscapes Encoder #Trainable 1/16 1/8 1/4 1/2
Networks (186) (372) (744) (1488)
Supervised Baseline R-50 $\times$1 63.3 70.2 73.1 76.6
PS-MT [30] [CVPR’22] R-50 $\times$1 - 75.8 76.9 77.6
U2PL [67] [CVPR’22] R-50 $\times$1 70.6 73.0 76.3 77.2
UniMatch [32] [CVPR’23] R-50 $\times$1 75.0 76.8 77.5 78.6
iMAS [40] [CVPR’23] R-50 $\times$1 74.3 77.4 78.1 79.3
AugSeg [39] [CVPR’23] R-50 $\times$1 73.7 76.5 78.8 79.3
CCVC [47] [CVPR’23] R-50 $\times$2 74.9 76.4 77.3 -
FPL (w/ CPS) [78] [CVPR’23] R-50 $\times$2 74.8 77.3 78.5 -
Diverse Co-T. (3-cps) [29] [ICCV’23] R-50 $\times$3 - 76.5 77.9 -
PrevMatch (ours) R-50 $\times$1 75.8 77.8 78.8 79.2
Supervised Baseline R-101 $\times$1 66.3 72.8 75.0 78.0
CPS [28] [CVPR’21] R-101 $\times$2 69.8 74.3 74.6 76.8
AEL [56] [NeurIPS’21] R-101 $\times$1 75.8 77.9 79.0 80.3
PS-MT [30] [CVPR’22] R-101 $\times$1 - 76.9 77.6 79.1
U2PL [67] [CVPR’22] R-101 $\times$1 74.9 76.5 78.5 79.1
PCR [57] [NeurIPS’22] R-101 $\times$1 73.4 76.3 78.4 79.1
CISC-R [79] [TPAMI’23] R-101 $\times$1 - 75.9 77.7 -
FPL (w/ AEL) [78] [CVPR’23] R-101 $\times$1 76.6 78.2 78.5 -
AugSeg [39] [CVPR’23] R-101 $\times$1 75.2 77.8 79.6 80.4
UniMatch [32] [CVPR’23] R-101 $\times$1 76.6 77.9 79.2 79.5
ESL [58] [ICCV’23] R-101 $\times$1 75.1 77.2 78.9 80.5
UPC (w/ U2PL) [80] [ICCV’23] R-101 $\times$1 75.3 77.4 79.0 79.6
Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 75.0 77.3 78.7 -
Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 75.7 77.4 78.5 -
Dual Teacher [31] [NeurIPS’23] R-101 $\times$1 76.8 78.4 79.5 80.5
PrevMatch (ours) R-101 $\times$1 77.7 78.9 80.1 80.1
@@
COCO & ADE20K.   Table V lists results for the large-scale datasets COCO and ADE20K. The proposed method consistently improves the performance across all partitions compared to the baseline, UniMatch, suggesting that PrevMatch is also effective on large-scale datasets.
TABLE V: Evaluation results on the large-scale datasets using Xception-65 for COCO and ResNet-50 for ADE20K.
Method COCO ADE20K
1/256 1/128 1/64 1/128 1/64 1/32
Supervised 28.0 33.6 37.8 7.2 9.9 13.7
UniMatch [32] 38.9 44.4 48.2 13.6 18.3 23.9
PrevMatch (ours) 40.2 45.7 48.4 15.4 19.6 24.9
4.3 Ablation Studies
In the ablation studies and discussion, the baseline refers to the UniMatch method [32].
4.3.1 Individual Efficacy of the Proposed Components
@@
The effects of the individual components are investigated, and the results are presented in Table VI. The first row indicates the baseline. All components of the proposed method consistently achieve performance gains. In particular, previous guidance, which randomly selects one previous model from the previous list for each iteration to generate additional guidance, surpasses the baseline by 0.8% and 1.3% in the 92- and 183-label settings, respectively. Regarding the number of models ($K$) for the network ensemble, we experiment with simple ensemble (fixed $K$) or random selection (random $K$) strategies. The result (third row) obtained using fixed $K$ exhibits minor performance gains. In contrast, the results using random $K$ (fourth row) indicate significantly improved performance. This implies that a fixed $K$ ensemble improves the reliability of pseudo-labels but limits the diversity, as noted in Dual Teacher [31], whereas a random $K$ strategy can provide reliable and diverse guidance to the model. In addition, using random weights for network ensembling contributes to performance gains, indicating that it expands the original pseudo-label space by generating diversified guidance.
TABLE VI: Ablation study of the components of the proposed method using a ResNet-50 encoder. For the ensemble, we set $K$=3. The means and standard deviation are also reported based on three runs to validate statistical significance.
Previous Simple Random Random PASCAL
Guidance Ensemble Selection Weights 92 183
- - - - 71.9±0.5 72.5±0.7
- - - 72.7±0.5 73.8±0.6
- - 72.7±0.4 74.1±0.5
- - 73.2±0.4 74.9±0.6
- 73.4±0.4 75.4±0.5
4.3.2 Previous List Length
@@
We investigate the effect of the length ($N$) of the previous list that stores the temporal models. In Table VII, the results for $N=1$ and $N=2$ are comparable to those of the baseline. This suggests that the aforementioned coupling problem may persist because the previous models in the list are continually updated with the latest model when $N$ is small. The cases of $N=$ 4, 8, and 12 consistently outperform the baseline, revealing that the proposed method is not highly sensitive to hyperparameters. However, the performance for $N=20$ increases only marginally due to the use of outdated teachers. Based on this result, we recommend setting the value of $N$ to approximately 5–15% of the total training epochs, which proves to be appropriate for different datasets (e.g., Pascal=6–10, Cityscapes=8–16, and COCO and ADE20k=4–5).
TABLE VII: Ablation study for the maximum length of the previous list using a ResNet-50 encoder. In this setting, only previous guidance is used (i.e., $K$=1, without ensemble).
List Length ($N$) Base. 1 2 4 8 12 20
@@
4.3.3 Upper Bound Number for Random Selection
To generate reliable and diverse pseudo-labels, we proposed a strategy that randomly selects $k$ models (ranging from 1 to $K$) for each iteration. In this strategy, we explore the performance changes regarding the upper bound number $K$. Table VIII indicates that including the ensembling cases ($K>1$, i.e., $k$=1 or $k>$1 are randomly selected) improves the performance significantly compared to the case of $K=1$ (i.e., without ensemble). In addition, we observe the best results at $K=3$ and a slight performance drop in settings with $K$ greater than 3. Even for large $K$, a varying number ($k$) of models is selected; however, the proportion of large $k$ values increases with $K$. This ensures consistent pseudo-labels but reduces their diversity, potentially degrading performance, as mentioned in Dual Teacher [31]. In conclusion, we select $K=3$ because it adequately satisfies the diversity and reliability requirements of the pseudo-labels.
TABLE VIII: Ablation study on the efficacy of the upper bound number $K$ using a ResNet-50 encoder and $N$=8.
@@ 4.3.3 Upper Bound Number for Random Selection
4.3.4 Criteria for Saving Previous Models
As described in Section 3.2, one alternative for storing previous models involves saving the model at regular intervals, a method used in [62, 63, 21]. Thus, we conduct experiments to validate the effectiveness of this approach. As listed in Table IX, although case (b) exhibits slightly better overall performance than the baseline (a), the difference is marginal. This suggests that storing models at short intervals does not address the coupling problem between the teacher and student networks. In contrast, case (c) shows a significant improvement compared to case (a). Although case (c) functions well, it exhibits limited improvements compared to case (d) which utilizes the proposed save criteria, demonstrating the superiority of the proposed approach. In addition, our approach does not require additional hyperparameter searches to determine appropriate intervals, thereby reducing unnecessary training costs.
TABLE IX: Ablation study regarding the efficacy of the save criteria using a ResNet-50 encoder, $N$=8, and $K$=3.
@@ 4.3.4 Criteria for Saving Previous Models
(d) On Best Epochs (PrevMatch) 73.4 75.4 77.5 78.6 79.3
Figure 3: Qualitative segmentation results on (a) Pascal VOC with a 92-label partition using a ResNet-50 encoder and (b) Cityscapes with a 1/16 label partition using a ResNet-101 encoder.
4.4 Discussion
TABLE X: Class-wise IoU scores for Pascal VOC with a 92-label partition using a ResNet-50 encoder. $\Delta$ indicates the difference between the PrevMatch and baseline UniMatch performances.

### v2

@@ 4.3 Comparison with State-of-the-Art Methods
PASCAL VOC. Two main experiments are conducted on the Original and Blended protocols. In Tab. 1, applying PrevMatch to each baseline consistently improves performance across all baseline methods. Notably, we observe that PrevMatch yields the most improvements in settings with fewer labels (92 and 183). In the 92-label setting (ResNet-101), it improves AugSeg, UniMatch, and MPMC by 2.3%, 1.8%, and 1.2%, respectively. Furthermore, even on UniMatchV2, which already exhibits strong performance with its transformer structure, PrevMatch further enhances performance. Moreover, Tab. 2 reports results on the Blended protocol. PrevMatch consistently improves the performance of all baseline methods.
Cityscapes. Experimental results for the Cityscapes validation set are presented in Tab. 3. The results indicate that applying PrevMatch to the four baseline approaches consistently improves their performance across all label partitions. Similar to the results reported in Tab. 1, PrevMatch significantly improves performance in setups with fewer labels. In particular, in the 1/16 label setup, PrevMatch improves AugSeg, UniMatch, UniMatchV2-S, and UniMatchV2-B by 1.4%, 1.1%, 1.1%, and 0.8%, respectively.
Cityscapes 1/16 1/8 1/4 1/2
(186) (372) (744) (1488)
ResNet-101
PS-MT [liu2022perturbed] - 76.9 77.6 79.1
Diverse Co-T.(3cps) [li2023diverse] 75.7 77.4 78.5 -
Dual Teacher [na2023switching] 76.8 78.4 79.5 80.5
CorrMatch [sun2024corrmatch] 77.3 78.5 79.4 80.4
AugSeg [zhao2023augmentation] 75.2 77.8 79.6 80.4
+ PrevMatch 76.6 78.9 80.3 80.8
Gain ($\Delta$) (+1.4) (+1.1) (+0.7) (+0.4)
UniMatch [yang2023revisiting] 76.6 77.9 79.2 79.5
+ PrevMatch 77.7 78.9 80.1 80.1
Gain ($\Delta$) (+1.1) (+1.0) (+0.9) (+0.6)
Tansformer-based
AllSpark [wang2024allspark] 78.3 79.2 80.6 81.4
SemiVL [hoyer2023semivl] 77.9 79.4 80.3 80.6
UniMatchV2-S [yang2025unimatch] 80.6 81.9 82.4 82.6
+ PrevMatch 81.7 82.7 83.2 83.2
Gain ($\Delta$) (+1.1) (+0.8) (+0.8) (+0.6)
UniMatchV2-B [yang2025unimatch] 82.6 83.3 83.5 83.9
+ PrevMatch 83.4 84.1 84.5 84.8
Gain ($\Delta$) (+0.8) (+0.8) (+1.0) (+0.9)
Table 3: Comparison with state-of-the-art methods on the Cityscapes dataset. $\dagger$: The results of UniMatchV2 (DINOv2-B) were reproduced with a training resolution of 7562 due to limitations in our available memory.
@@
ADE20K. Tab. 4 lists results for the large-scale and challenging dataset ADE20K. PrevMatch consistently improves baseline performance across all partitions, suggesting that PrevMatch is also effective on the large-scale dataset.
ADE20K 1/128 1/64 1/32 1/16 1/8
UniMatch-RN50 [yang2023revisiting] 13.6 18.3 23.9 27.2 30.9
+ PrevMatch 15.4 19.6 24.9 28.3 31.6
Gain ($\Delta$) (+1.8) (+1.3) (+1.0) (+1.1) (+0.7)
UniMatchV2-S [yang2025unimatch] 20.9 29.1 34.5 37.0 39.4
+ PrevMatch 22.2 29.8 35.2 37.8 40.0
Gain ($\Delta$) (+1.3) (+0.7) (+0.7) (+0.8) (+0.6)
UniMatchV2-B [yang2025unimatch] 25.0 33.7 40.5 42.4 45.1
+ PrevMatch 25.7 35.1 41.4 43.4 46.0
Gain ($\Delta$) (+0.7) (+1.4) (+0.9) (+1.0) (+0.9)
Table 4: Evaluation results on the large-scale dataset ADE20K, as reproduced in our experimental setup.
Previous Simple Random Random w/ UniM. w/ UniM.V2
Guidance Ensemble Selection Weights 92 92
- - - - 71.9 79.0
- - - 72.7 82.2
- - 72.7 82.7
- - 73.2 83.8
- 73.4 84.4
Table 5: Ablation study of PrevMatch’s components based on UniMatch (ResNet-50) and UniMatch V2 (DINOv2-S) on Pascal VOC with 92 labels. For the ensemble, we set KK=3.
5 Discussion
5.1 Ablation Study
@@
The effects of the individual components are investigated, and the results are presented in Tab. 5. The first row indicates the baselines. All components of the proposed method consistently achieve performance gains. In particular, previous guidance, which randomly selects one previous model from the previous list for each iteration to generate additional guidance, surpasses the baselines by 0.8% and 3.2% in the 92-label settings, respectively. Regarding the number of models ($k$) for the network ensemble, we experiment with simple ensemble (fixed $k$) or random selection (random $k$) strategies. The result (third row) obtained using fixed $k$ exhibits minor performance gains. In contrast, the results using random $k$ (fourth row) indicate significantly improved performance. This implies that a fixed $k$ ensemble improves the reliability of pseudo-labels but limits the diversity, as noted in Dual Teacher [na2023switching], whereas a random $k$ strategy can provide reliable and diverse guidance to the model. In addition, using random weights for network ensembling contributes to performance gains, indicating that it expands the original pseudo-label space by generating diversified guidance. Refer to Tabs. 10, 11, 13 and 12 in the supplementary material for additional ablation studies on different protocols and the hyperparameters of each component.
chair
sofa
car
train
m.bike
bicycle
d.table
backgr.
UniMatch 8 33 80 80 75 59 55 91
+ PrevMatch 21 46 84 84 77 61 57 93
Gain ($\Delta$) $cellcolorgreen!3013 \cellcolorgreen$$\cellcolorgreen!154$ $cellcolorgreen!102 \cellcolorgreen$$\cellcolorgreen!102$
Table 6: Class-wise IoU scores (Top 8 classes based on gain) for Pascal VOC with a 92-label partition using ResNet-50.
5.2 Analysis of Performance and Training Stability
Tab. 6 lists the category-wise IoU scores and shows that the proposed method achieves notable performance gains for the chair and sofa classes, which were particularly challenging for the UniMatch baseline. To further explore these poorly behaved classes, we visualize changes in terms of pseudo-label accuracy and validation IoU scores during training, as depicted in Fig. 2. In the first row (chair class), the training curve of the baseline pseudo-label accuracy exhibits significant fluctuations, particularly showing a sudden sharp performance drop at approximately 50 epochs. Although the pseudo-label accuracy recovers slightly thereafter, the validation score does not. In the second row (sofa class), the training curve of the baseline exhibits more severe fluctuations and sharper and more drastic performance drops than in the first row. This phenomenon can be attributed to catastrophic forgetting where previously learned knowledge is forgotten when acquiring new knowledge [mccloskey1989catastrophic, kirkpatrick2017overcoming]. In other words, self-training and class-imbalance scenarios (i.e., in our scenario) lead to instability in pseudo-label predictions, particularly for poorly behaved classes, due to the lack of labeled data. In contrast to the baseline, the proposed method shows a smoother training curve without significant fluctuations in either category. This indicates that the proposed training procedure aids in achieving stable optimization for poorly behaved classes that suffer from the forgetting problem (Refer to Sec. 8 in the supplementary material for additional analysis).
Figure 2: Training curves for the chair and sofa classes, illustrating variations in pseudo-label pixel accuracy and validation IoU scores (on Pascal VOC 92-label partition).
Method Encoder Pascal VOC $\rightarrow$ COCO
UniMatch [yang2023revisiting] RN-50 71.9 35.6
+ PrevMatch RN-50 73.4 (+1.5) 38.5 (+2.9)
UniMatchV2-B [yang2023revisiting] DINOv2-B 86.3 52.9
+ PrevMatch DINOv2-B 87.2 (+0.9) 55.1 (+2.2)
Table 7: Evaluation results on seen (Pascal VOC) and unseen (COCO) datasets. COCO contains 80 object classes, 20 of which correspond to the categories in Pascal VOC. The remaining 60 classes were treated as a background class for evaluation. The results on COCO were obtained without re-training, using the weights of the model trained on Pascal VOC with a 92-label setting.
5.3 Training Stability and Generalization
Several studies [li2018visualizing, foret2020sharpness, chen2021vision, shin2023metricgan] across various domains have demonstrated that stable optimization during training results in a flatter loss landscape, thereby improving generalization performance. Therefore, to assess whether the proposed method’s training stability contributes to improved generalization, we conduct additional experiments. Tab. 7 presents evaluation results for models trained on Pascal VOC [everingham2010pascal] and tested on the COCO dataset [lin2014microsoft] without re-training. The results show a larger difference in the score on the unseen dataset (COCO) as compared to that on the seen dataset (Pascal VOC), indicating that PrevMatch improves the model’s generalization ability.
Method #Trainable Networks Training Time (m) GPU Mem. Usage (G) Pascal
92 183 366
PS-MT 1 24.9 28.6 65.8 69.6 76.6
Diverse Co-T. 3 22.1 40.5 75.4 76.8 79.6
UniMatch 1 7.8 21.9 75.2 77.2 78.8
+ PrevMatch 1 9.1 22.7 77.0 78.5 79.6
Table 8: Comparison of time spent using existing methods with a ResNet-101 encoder. Training time per epoch and GPU memory usage were measured using the same environment (two A6000 GPUs) and hyperparameters. We used the open-sourced code provided by the authors.
5.4 Efficiency Evaluation
To evaluate the efficiency of PrevMatch, we compare its training time per epoch and GPU memory usage with existing methods. Since UniMatch [yang2023revisiting] relies solely on input separation, it may suffer from instability due to the inherent coupling problem. In contrast, PS-MT [liu2022perturbed], Diverse Co-training [li2023diverse], and our proposed PrevMatch address this limitation by integrating both network and input separation techniques. In Tab. 8, PrevMatch improves UniMatch by 1.8% (92-label setting), with a slight cost increase. Moreover, PrevMatch outperforms Diverse Co-training, while achieving 2.4 times faster training. Meanwhile, PS-MT, a dual Mean Teacher-based method, requires training time comparable to that of Diverse Co-training due to its complex components, despite being a single trainable network. These results suggest that PrevMatch can efficiently mitigate the coupling problem. Therefore, owing to its computational efficiency and the simplicity of its pipeline reusing the inputs and networks, the proposed method can be easily integrated into any semi-supervised learning method.
6 Conclusion
In this work, we introduced the PrevMatch framework, which leverages temporal knowledge obtained during training to efficiently address the issues of tight coupling and confirmation bias that impede stable semi-supervised learning. The main contributions of PrevMatch include revisiting the use of temporal knowledge and maximizing its effectiveness. Experiments were conducted on three benchmark datasets, revealing that incorporating PrevMatch into existing methods significantly improves their performance. In addition, our findings indicate that leveraging temporal knowledge facilitates stable optimization, particularly for classes that exhibit poor performance and fluctuations. Finally, the computational efficiency and compatibility of the proposed method facilitate its seamless integration into recent semi-supervised semantic segmentation methods.
Acknowledgements.   This research was supported by the BK21 FOUR funded by the Ministry of Education of Korea and National Research Foundation of Korea.
[…]
Supplementary Material
7 Additional Experiments and Ablation Studies
In this study, experiments were conducted in the following environments: Ubuntu 20.04, Python 3.10.4, PyTorch 1.12.1, and NVIDIA 3090Ti or A6000 GPUs. Unless otherwise specified, subsequent experiments were primarily conducted using the UniMatch baseline with ResNet-50.
7.1 Comparison with State-of-the-Art on the Priority Protocol of Pascal VOC
As described in the main paper, the Pascal VOC dataset contains a high-quality annotation subset (1,464 images) and a coarse-quality annotation subset (9,118 images). In addition to the Original and Blended protocols presented in the main manuscript, the Priority protocol has been used in several studies for comparison. The protocol where the selection of labeled images is first derived from the high-quality subset; if not sufficient, it is complemented by additional images from the coarse subset. Tab. 9 shows that the proposed method consistently improves baseline methods.
7.2 Additional Ablation Studies
Due to space constraints, the ablation study (Tab. 5) in the main body of the paper includes only the 92-label protocol. To further validate the proposed method, we provide additional experiments on the 183- to 1464-label protocols. Tab. 10 shows that the components of PrevMatch consistently improve performance across various evaluation protocols.
7.3 Previous List Length
@@
We investigate the effect of the length ($N$) of the previous list that stores the temporal models. In Tab. 11, the results for $N=1$ and $N=2$ are comparable to those of the baseline. This suggests that the aforementioned coupling problem may persist because the previous models in the list are continually updated with the latest model when $N$ is small. The cases of $N=$ 4, 8, and 12 consistently outperform the baseline, revealing that the proposed method is not highly sensitive to hyperparameters. However, the performance for $N=20$ increases only marginally due to the use of outdated teachers. Based on this result, we recommend setting the value of $N$ to approximately 5–15% of the total training epochs, which proves to be appropriate for different datasets (e.g., Pascal=6–10, Cityscapes=8–16, and ADE20k=4–6).
Pascal [Priority set] Encoder 1/16 1/8 1/4
U2PL [wang2022semi] RN-101 77.2 79.0 79.3
U2PL+ [wang2024using] RN-101 77.2 79.4 80.2
Dual Teacher [na2023switching] RN-101 80.1 81.5 80.5
CorrMatch [sun2024corrmatch] RN-101 81.3 81.9 80.9
AllSpark [wang2024allspark] MiT-B5 81.6 82.0 80.9
FixMatch [sohn2020fixmatch] RN-50 75.4 76.5 76.9
+ PrevMatch RN-50 76.9 77.6 77.8
     Gain ($\Delta$) (+1.5) (+1.1) (+0.9)
UniMatch [yang2023revisiting] RN-101 80.9 81.9 80.4
+ PrevMatch RN-101 81.4 81.9 80.8
     Gain ($\Delta$) (+0.5) (+0.0) (+0.4)
UniMatchV2-S [yang2025unimatch] DINOv2-S 86.6 87.3 85.9
+ PrevMatch DINOv2-S 87.3 87.8 86.4
     Gain ($\Delta$) (+0.7) (+0.5) (+0.5)
Table 9: Comparison with state-of-the-art methods on the Priority protocol of Pascal VOC dataset.
Previous Random Random PASCAL
Guidance Selection Weights 92 183 366 732 1464
- - - 71.9 72.5 76.0 77.4 78.7
- - 72.7 73.8 76.6 78.1 79.0
- 73.2 74.9 77.4 78.3 79.2
73.4 75.4 77.5 78.6 79.3
Table 10: Additional ablation studies across different evaluation protocols (UniMatch with ResNet-50).
NN Base. 1 2 4 8 12 20
@@
Table 11: Ablation study for the maximum length ($N$) of the previous list. In this setting, only previous guidance is used (i.e., KK=1, without ensemble).
7.4 Upper Bound Number for Random Selection
To generate reliable and diverse pseudo-labels, we proposed a strategy that randomly selects $k$ models (ranging from 1 to $K$) for each iteration. In this strategy, we explore the performance changes regarding the upper bound number $K$. Tab. 12 indicates that including the ensembling cases ($K>1$, i.e., kk=1 or k>k>1 are randomly selected) improves the performance significantly compared to the case of $K=1$ (i.e., without ensemble). In addition, we observe the best results at $K=3$ and a slight performance drop in settings with $K$ greater than 3. Even for large $K$, a varying number ($k$) of models is selected; however, the proportion of large $k$ values increases with $K$. This ensures consistent pseudo-labels but reduces their diversity, potentially degrading performance, as mentioned in Dual Teacher. In conclusion, we select $K=3$ because it adequately satisfies the diversity and reliability requirements of the pseudo-labels.
@@ 7.4 Upper Bound Number for Random Selection
Table 12: Ablation study on the efficacy of the upper bound number $K$ using NN=8.
@@
(d) On Best Epochs (ours) 73.4 75.4 77.5 78.6 79.3
Table 13: Ablation study regarding the efficacy of the save criteria using NN=8 and KK=3.
7.5 Criteria for Saving Previous Models
As described in Sec. 3.2.1 of the main paper, one alternative for storing previous models involves saving the model at regular intervals, a method used in [huang2017snapshot, wang2022efficient, yang2022st++]. Thus, we conduct experiments to validate the effectiveness of this approach. As listed in Tab. 13, although case (b) exhibits slightly better overall performance than the baseline (a), the difference is marginal. This suggests that storing models at short intervals does not address the coupling problem between the teacher and student networks. In contrast, case (c) shows a significant improvement compared to case (a). Although case (c) functions well, it exhibits limited improvements compared to case (d) which utilizes the proposed save criteria, demonstrating the superiority of the proposed approach. In addition, our approach does not require additional hyperparameter searches to determine appropriate intervals, thereby reducing unnecessary training costs.
7.6 Loss Weight
The loss weight $\lambda$ for previous guidance plays a critical role in the overall training process. Since the primary goal of previous guidance is to mitigate confirmation bias–which tends to occur during the early to middle stages of training–its effect is most beneficial before the model becomes overconfident in incorrect predictions. However, in the early phase, the model’s predictions are typically unreliable. To address this, we employ a warmup schedule that gradually increases $\lambda$, similar to commonly used learning rate warmup strategies. Furthermore, as training progresses and the model becomes more stable and accurate, the need for strong regularization from previous guidance naturally decreases. To reflect this, we apply a decay to $\lambda$ in the later stages. Our final schedule (row (e)) allows previous guidance to act most strongly when it is most needed–during the middle of training–while reducing its influence as the model converges.
Tab. 14 presents an ablation study comparing different $\lambda$ scheduling strategies. The fixed (b) and linear decay (c) settings maintain a high weight even in the early stage, where the model’s predictions are still unstable. As a result, these configurations show limited improvements, likely due to the guidance based on noisy predictions. In contrast, the linear increase (d) and warmup+decay (e) strategies both incorporate a warmup phase, mitigating this issue and yielding significant performance gains across various label partitions. While both (d) and (e) outperform other variants, the warmup+decay strategy shows slightly better or comparable results, especially in high-label settings. This indicates that reducing the influence of previous guidance in the late stage, when the model is already well-trained, is beneficial for final performance.
Loss weight ($\lambda$) 92 183 366 732 1464
(a) Baseline 71.9 72.5 76.0 77.4 78.7
(b) Fixed (1.0) 72.1 74.2 76.7 78.2 78.6
[…]
Table 14: Ablation study for $\lambda$.
8 Additional Analysis
Figure 3: Qualitative segmentation results on (a) Pascal VOC and (b) Cityscapes.